Evening Shift 11/18/99

Peter Shanahan (shanahan@fnal.gov)
Fri, 19 Nov 1999 00:01:00 -0600

Shift Summary for eve shift, 18th Nov 99

On Shift: PSH, PMcB

Hours of beam available at all intensities: 8.25
Hours of beam available at nominal intensity: 6.00
Hours of E799 data (special runs not included): 8.15
Hours of special runs (laser scans, muon runs...): 0.00

Accounting of lost time:
H-Source problems (15')
Start/stop of runs
KRENDS

========================================================================
Runs ended this shift:

Run 15004 - 12.2M events - BEAM_799
Run 15005 - useless - COSMIC
Run 15006 - 8.0M events - BEAM_799

Tapes brought to FCC this shift:
(including tapes for online split)

none

========================================================================
Accesses during shift:

none

========================================================================
Overview of shift:
E799 Data Taking

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------

15004 260 16:50 Called MCR to complain about dropping intensity.
They say it's being looked into.
15004 17:25 DN switches to He tubes 29&30.
15004 18:15 Page DN about CO_2 pressure. He says to read
gauge connected directly to Dewar, as opposed to
regulator gauge.
15004 18:27 H-Source problems, no beam. Page Ashkan to
start new SDAQ DB. We stop run, to take cosmic
latch data.
15005 18:40 COSMIC run - no good data. Many kfrends/pipe inits.
18:45 Beam returns. Stop run, init BEAM_799
15006 1 18:53 Start BEAM_799 run.
19:05 Page Jason H. about region of HCC with ~0.5% ineff.
He says it was probably Etotal problem that went
away with new run initialization.
20:15 Mount KQH026 on ktev2.
21:20 1 filter process on ktev1 is taking ~30% more CPU
than others, and is just finishing in time. Event
throughput is flat. Leo says to keep going, but
fresh start or force level-3 on next init.
15006 22:45 Run ends due to full tapes, but DAQMON says only
80% full. Sydney investigates. Fresh start.
15007 1 23:01 Start run. KFREND, Load KQs, KFREND.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

Take a 15' cosmic run for VVH latch studies, if beam goes away for
more than a few minutes.

========================================================================
Items for Experts and Management

o RC ADC distribution has a few events over limit in kumac. It
was mentioned several days ago, but doesn't look serious. Update
limits?

o Is l3#cpu kumac out of date with respect to new l3 config?

o VVH latch timing still a dog.

o montup calculation of ET efficiency uses too low a cut for ET4,
giving artificial inefficiencies.

o Sydney believes tape compression ratio in DB might be incorrect.

=======================================================================
Fermilab MS-231 Phone: (+1) 630 840 8378
P.O. Box 500 Fax: (+1) 630 840 6311
Batavia, IL 60510 USA